OpenStudy (sidmanfu):

If the base of a triangle is 62 m and the area is 186 m², what is the height of the triangle?

OpenStudy (anonymous):

\[ \frac 12\times h \times 62 =186 \] can you solve for h?

OpenStudy (anonymous):

h = 6.

OpenStudy (anonymous):

Do you understand?

OpenStudy (sidmanfu):

How did u get that answer? I still don't get it.

OpenStudy (anonymous):

Area of a triangle is given by \[\frac 12 \times Base \times height \]

OpenStudy (sidmanfu):

yes i know that.